I need to know how to multiply (5- square root of 11) by it's conjugate and simplify. Help please?
Answer by stanbon(48568)

I need to know how to multiply (5- square root of 11) by it's conjugate and simplify.
----------------------
Conjugates multiplication takes the following form:
(a - sqrt(b))(a + sqrt(b)) = a^2 -b
---------------------------------------------
(5- sqrt(11))(5 + sqrt(11)) = 25 - 11 = 14
===============================================
Cheers,
Stan H>